# FAIR Universe HiggsML Uncertainty Challenge Competition

### v3

@@ 5 Baselines, code, and material provided
The Baseline method estimates $\mu$ by using standard techniques without addressing systematic uncertainties directly for simplicity. Initially, it utilizes a classifier (based on XGBoost Boosted Decision Tree, or a simple PyTorch dense Neural Network) trained on BaselineTrain, a subset of training data to enhance signal event density and reduce $\mu$ estimator variance. Although adjustable for variance optimization, the classifier’s decision threshold is fixed heuristically. $\mu$ is then estimated from these filtered events, assuming Poisson distribution, enabling point-wise and interval maximum likelihood estimation. Further refinement involves binning particles as per classifier score and estimating $\mu$ per bin, akin to a voting ensemble. BaselineHoldout, a reserved training dataset, is used to predict the amount of background and signal in each bin for $\mu$=1. This calibration step then permits estimating $\mu$ (and the corresponding Confidence Interval) on pseudo-experiment for $\mu$ estimation. On Figure 4, the alignment of maximum likelihood estimation (orange line) with unlabelled data (black line) indicates method success, in the absence of any bias.
@@
Now, when unknown biases occur, the prediction on the amount of background and signal event per bin will be wrong, biasing the estimation of $\mu$. To address the problem of systematic errors, participants are encouraged to enhance the Baseline model, for instance, by modifying the architecture or training protocol to improve resilience against biases, attempting to directly model the biases, or refining the estimator through a bias-aware model. Another way to see it is that, armed with the biasing script which can produce a dataset for any value of the six Nuisance Parameters and the signal strength $\mu$, the participants could train a model which could regress the seven parameters for any pseudo-experiment and report the Confidence Interval on $\mu$.
@@
Figure 4: Model Output: Unlabelled test pseudo-data (black points), and HoldOut data for (1) background events $Z\rightarrow\tau\tau$ (solid green), (2) background $t\bar{t}$ (solid orange) (3) background di-boson (solid blue, hardly visible) (4) signal events $H\rightarrow\tau\tau$ for $\mu=1$ (red line), and (5) signal events fitted histogram to test pseudo-data, leading to estimated $\mu=2.667$ (orange line).
5.1 Website, tutorial and documentation
We have set up a dedicated GitHub repository [^2] https://github.com/FAIR-Universe/HEP-Challenge for this competition and the competition’s Codabench website [^3] https://www.codabench.org/competitions/2977/. The GitHub repository covers the following points:
  • Competition introduction and instructions for setting up the complete environment.
  • Complete details of the evaluation process.
  • Information about how to submit.
  • Troubleshooting instructions for possible issues and contact details for reporting issues.
  • Link to a dedicated forum on the Codabench platform for efficient communication.
In addition, a code tutorial is provided for: (i) loading and discovering properties of data. (ii) explaining the coding structure and expected functions to be implemented in submissions. (iii) providing instructions and examples for running the baseline methods on the public datasets.
6 Organizational aspects
6.1 Overview
Competition
The competition is hosted on Codabench platform with a dedicated webpage on Codabench. Participants are required to create accounts on the platform and register for the competition, accepting the competition’s rules, as listed on the site Participants can see and access all the public details without registering for the competition. Both signing up to Codabench and registering for the competition have no fee.
Competition Phases
The competition consists of two phases:
  • Phase 1: Public Phase – Participants submit their model
  • Phase 2: Final Phase – Participants now submit only one best submission from the previous phase for final evaluation (by default, it would be one with the highest score).
Competition Pages
Participants are provided with multiple pages detailing the competition overview, evaluation procedures, data description, competition terms, starting kit, and public data.
Starting Kit
The starting kit is a tutorial notebook designed to familiarise participants with the competition protocol, competition data, scoring, and code to prepare their submissions for the competition website. Participants are also given some ready-to-submit submissions in the GitHub repository to get them started with libraries such as PyTorch pytorch , TensorFlow tensorflow , scikit-learn scikit-learn etc.
Submissions
Once registered, participants can submit their solutions (code submissions, possibly including training), which are executed on resources at the National Energy Research Scientific Computing Center (NERSC). Participants receive submission feedback once per day, as displayed on the competition leaderboard. Every submission is executed in the same environment with the same maximum execution time of 2 hours (for the complete evaluation), translating into about 10 second GPU maximum inference time for one pseudo-experiment. Each participant is allowed to make only five submissions per day and a maximum of 100 submissions during the challenge. To enable the participants to perform other experiments on their hardware, they can use the public data provided in Phase 1.
Codabench Tutorial
We have prepared a codabench tutorial for participants to get started with the competition quickly. The tutorial gives all the steps to access the competition, public data, starting kit, dummy submission, and leaderboard. Furthermore, we have detailed the competition flow, submission process, and contact information at the end of the tutorial.
6.2 Competition Protocol
Participants can either submit a pre-trained model or train their model as part of the submission. When a participant uploads the submission to the competition, that code is executed by the following two steps (summarized in Figure 5 and Figure 6).
Ingestion
The ingestion step executes an ingestion program that is responsible for the following:
  • loads train data
  • loads test data (without labels)
  • initializes the submitted model
  • calls the fit method of the submitted model (to be used for training a model or loading a pre-trained model that is submitted together with the submission)
  • calls the predict method multiple times (once for each test set) to get predictions (mu_hat, delta_mu_hat, p16, and p84)
  • saves the predictions for the scoring program
Figure 5: Ingestion Program execution flow
Scoring
The scoring step executes a scoring program that is responsible for the following:
  • loads the predictions saved by the ingestion program
  • loads the ground truth or test settings
  • calculates scores
  • saves scores and detailed results
Figure 6: Scoring Program execution flow
@@ 7 Conclusions and Outlook
We have prepared a dataset, challenge, and platform for developing and comparing machine learning methods that quantify uncertainties in addition to providing point estimates. This project builds on previous work in the context of the Higgs boson discovery at the Large Hadron Collider. With the growing size of datasets in high energy physics, the sophistication of tools, and the precision requirements to explore new phenomena, uncertainty quantification will be an essential part of machine learning going forward.
Building a challenge of this kind has itself been a challenge. We needed to ensure that the problem was realistic enough, manageable to non-experts, and useful. These requirements were often in conflict, e.g. we wanted systematic uncertainties to be most important, but we also needed the dataset size to not be too large so that non-experts could analyse them. We also appreciate that there is no unique metric for trading off accuracy and precision and we look forward to further discussions on the choice we made throughout the challenge period and beyond [^4] The dataset will continue to be available after the challenge and we hope that it will continue to provide a benchmark for method developments in the future..
While focused on high energy physics and the Higgs boson, the features of the competition are much more general. We hope that the lessons learned here will be valuable across high energy physics and beyond.

### v4

@@ 4 Software
The Starting Kit kit includes code for installing necessary packages, loading and visualising data, training and evaluating a model with the metrics described in subsection 3.1. The Baseline method estimates $\mu$ using standard techniques without directly addressing systematic uncertainties for simplicity. Initially, it utilises a classifier (based on an XGBoost Boosted Decision Tree) trained on a subset of the training data to enhance the signal event density and reduce the $\mu$ estimator variance. The classifier’s decision threshold is fixed heuristically. $\mu$ is then estimated from these filtered events, assuming a Poisson distribution, enabling interval maximum likelihood estimation. Further refinement involves binning events based on their classifier score and estimating $\mu$ per bin. A holdout dataset, is used to predict the amount of background and signal in each bin for $\mu=1$. This calibration step then permits estimating $\mu$ (and the corresponding CI) on each pseudo-experiment. On 2(a), the alignment of maximum likelihood estimation (orange line) with unlabelled data (black line) indicates the method success, in the absence of any bias.
@@
When unknown biases occur, the prediction on the amount of background and signal events per bin will be wrong, biasing the estimation of $\mu$. To address the problem of systematic errors, we use the holdout dataset with biases by different amounts of the Nuisance Parameter ($\theta$) and then build a calibration curve to estimate the signal and background in each bin. 2(b) shows one such fit curve for the 24th bin (just as an example). Now, instead of $\mu$ depending on $S$ and $B$, it will depend on fit functions $S(\theta)$ and $B(\theta)$. Finally, the minimisation function now regresses both $\mu$ and $\theta$, thus making the model less susceptible to systematic bias. But this is only limited to one nuisance parameter; participants are encouraged to enhance the Baseline model, for instance, by modifying the architecture or training protocol to improve resilience against biases, attempting to directly model the biases, or refining the estimator through a bias-aware model.
Another way to see it is that, armed with the biasing script which can produce a dataset for any value of the six Nuisance Parameters and the signal strength $\mu$, the participants could train a model which could regress the seven parameters for any pseudo-experiment and report the Confidence Interval on $\mu$. This was actually done with different techniques by the winning trio (section 5).
@@
Figure 2: (a) classifier score for unlabelled test data (black points), and holdout data for (1) background events $Z\rightarrow\tau\tau$ (solid green), (2) background $t\bar{t}$ (solid orange) (3) background di-boson (solid blue, hardly visible) (4) signal events $H\rightarrow\tau\tau$ for $\mu=1$ (red line), and (5) signal events fitted histogram to test data, leading to estimated $\mu=2.667$ (orange line) (b) model of the bin content vs Nuisance Parameter $\theta$ for bin 24, as an example.
5 Competition results and best submissions
At the end of the competition, a clear trio was at the top of the public leaderboard: HEPHY with a quantile score of 0.878, followed by Ibrahime (0.823) and Hzume (0.179). All submissions have been reevaluated on a new dataset (i.i.d. to the original one). The evaluation was done on 1000 trials of 100 pseudo-experiments (each trial with a given value of $\mu$ randomised between 0.1 and 3), instead of 10 trials for the public leaderboard. All submissions were run on the same pseudo-experiments, instead of separate pseudo-experiments for the public leaderboard.
Figure 3 shows the results for all trials for the trio. The CI width is seen falling at large values of $\mu$: this is due to the clipping the Confidence Interval to a maximum value of 3 (which was not done in 1(a)), which was the maximum value in this competition. Such clipping would be meaningless in the context of a real physics measurement where $\mu$ is truly unknown. This is the only "hack" specific to the competition context that could be identified. As far as the score is concerned, HEPHY and Ibrahim are very close. When merging all trials, the scores obtained by the top trio are: HEPHY -0.582, Ibrahim -0.576 and HZUME -2.16. An additional bootstrap analysis of the variance of these results showed that HEPHY and Ibrahim cannot be reliably ranked, hence the final rankings :
Figure 3: Comparative study of the three finalists (blue for Hzume, orange for HEPHY and green for Ibrahim’s model) with 1000 trials of 100 pseudo-experiments (see subsection 3.1). 3(a) the coverage from each trial, 3(b) the average CI width and 3(c) the quantile score
  • 1st tie: HEPHY (Lisa Benato, Cristina Giordano, Claudius Krause, Ang Li,  Robert Schöfbeck, Maryam Shooshtari, Dennis Schwarz, Daohan Wang) from Vienna’s Institute of High Energy Physics (HEPHY) in Austria wins \$2000.
  • 1st tie IBRAHIME (Ibrahim Elsharkawy) from University of Illinois at Urbana-Champaign, USA wins \$2000.
  • 3rd HZUME (Hashizume Yota) from Kyoto University, Japan wins \$500
All three are co-authors of this paper and have provided a summary of their algorithms in the following sub-sections. HEPHY and Ibrahim’s sub-sections also refer to their public full papers and their code.
5.1 HEPHY: Simulation-based inference with a calibrated multiclassifier and parametric regressors for learning systematics
[…]
A multiclass classifier is trained on nominal (i.e., unvaried) simulation data and predicts the class probabilities for the four processes: $\mathrm{H}\to\tau\tau$, $\mathrm{Z}\to\tau\tau$, $\mathrm{t}\overline{\mathrm{t}}$, and $\mathrm{VV}$. The output class probability is scaled with $(1+\alpha)^{\nu}$ for each of the nuisance parameters $\nu_{\text{bkg}}$, $\nu_{\mathrm{t}\overline{\mathrm{t}}}$, and $\nu_{\mathrm{VV}}$, that control the normalization of background processes. The three accompanying constants $\alpha$ determine the pre-fit sizes of these uncertainties. A critical step is a dedicated and highly precise iterative isotonic regression step to calibrate the classifier’s output.
To account for the dependence of the likelihood on the remaining systematic uncertainties, a second set of networks estimates the relative variation in the differential cross section as a function of ${\boldsymbol{\nu}}_{\text{calib}}=\{\nu_{\text{tes}},\nu_{\text{jes}},\nu_{\text{met}}\}$. These nuisances control uncertainties in the calibration of the data and enter training data via the biasing script. We fit an exponential ansatz parameterised by a neural network for each of the four processes (labeled by $p$) and separately in each region: $\frac{{\textrm{d}}\sigma_{p}(\boldsymbol{x}|\mu,{\boldsymbol{\nu}})}{{\textrm{d}}\sigma_{p}(\boldsymbol{x}|1,{\mathbf{0}})}\simeq\hat{S}_{p}(\boldsymbol{x}|{\boldsymbol{\nu}}_{\text{calib}})=\exp(\nu_{A}\hat{\Delta}_{r,p,A}(\boldsymbol{x}))$, where $\nu_{A}$ is a multi-index that labels three linear, three quadratic, and three mixed terms of the three calibration-type nuisances. The $\hat{\Delta}_{r,p,A}(\boldsymbol{x})$ are functions learned by the network and specific to the selection $r$ and the process $p$. Based on the cross-entropy loss, the ansatz leads to the loss function
$L[\hat{\Delta}_{A}]=\sum_{{\boldsymbol{\nu}}\in\mathcal{V}}\Bigg[\int{\textrm{d}}\sigma(\boldsymbol{x}|{\mathbf{0}})\,\textrm{Soft}^{+}(\nu_{A}\hat{\Delta}_{A}(\boldsymbol{x}))+\int{\textrm{d}}\sigma(\boldsymbol{x}|{\boldsymbol{\nu}})\,\textrm{Soft}^{+}(-\nu_{A}\hat{\Delta}_{A}(\boldsymbol{x}))\Bigg].$ (3)
This architecture allows the surrogate to interpolate continuously in both feature and nuisance parameter space. The complete likelihood can then be computed from the surrogate for the differential cross-section ratio with the closed-form expression
$\frac{{\textrm{d}}\sigma(\boldsymbol{x}|\mu,{\boldsymbol{\nu}})}{{\textrm{d}}\sigma(\boldsymbol{x}|1,{\mathbf{0}})\,}$ $\simeq\mu\hat{g}_{\mathrm{H}}(\boldsymbol{x})\hat{S}_{\mathrm{H}}(\boldsymbol{x}|{\boldsymbol{\nu}}_{\text{calib}})+(1+\alpha_{\text{bkg}})^{\nu_{\text{bkg}}}\Big(\hat{g}_{\mathrm{Z}}(\boldsymbol{x})\hat{S}_{\mathrm{Z}}(\boldsymbol{x}|{\boldsymbol{\nu}}_{\text{calib}})$
$+(1+\alpha_{\mathrm{t}\overline{\mathrm{t}}})^{\nu_{\mathrm{t}\overline{\mathrm{t}}}}\,\hat{g}_{\mathrm{t}\overline{\mathrm{t}}}(\boldsymbol{x})\hat{S}_{\mathrm{t}\overline{\mathrm{t}}}(\boldsymbol{x}|{\boldsymbol{\nu}}_{\text{calib}})+(1+\alpha_{\mathrm{VV}})^{\nu_{\mathrm{VV}}}\,\hat{g}_{\mathrm{VV}}(\boldsymbol{x})\hat{S}_{\mathrm{VV}}(\boldsymbol{x}|{\boldsymbol{\nu}}_{\text{calib}})\Big)$ (4)
where $\hat{g}_{p}(\boldsymbol{x})$ is the (calibrated) output of the multiclassifier. The surrogate is efficient in evaluating and differentiable with respect to all parameters. For the inclusive cross-section component of the extended likelihood, we introduce a spline-based interpolation scheme that reduces numerical instabilities and speeds up the evaluation during profiling.
We train one multiclass classifier and one systematic network per selection. Closure tests show that the surrogates reproduce the shapes and normalisations of the simulated distributions across many kinematic observables and several orders of magnitude. The unbinned surrogate could be further refined in a modular way: new systematics or background processes can be added without retraining the entire model, mirroring the workflow of traditional HEP data analyses. This “refinable” modelling is crucial for scalability in real LHC analyses where hundreds of nuisance parameters are typical.
Figure 4: Scatter plot of the true value of the $\mathrm{H}\to\tau\tau$ signal strength parameter $\mu$ (left) and the MLE $\bar{\mu}$ for $5\cdot 10^{4}$ toys showing stability over the whole range of relevant $\mu_{\text{true}}$. The normalisation-type nuisance parameter $\nu_{\text{$\mathrm{t}\overline{\mathrm{t}}$}}$ (middle) and the calibration-type nuisance parameter $\nu_{\text{jes}}$ (right) are severely constrained, reducing the impact of the corresponding uncertainties.
We profile the nuisance parameters using the MINUIT package [78] and determine the 68% CI by evaluating the profiled likelihood as a function of $\mu$. The gain from the unbinned model becomes evident when performing inference, where the unbinned surrogate model improves the expected $1\sigma$ confidence interval on the signal strength by 20% compared to a traditional binned analysis using classifier-based templates. The unbinned model also leads to significantly stronger constraints on nuisance parameters, especially for calibration-related systematics like $\nu_{\text{tes}}$ and $\nu_{\text{jes}}$, reducing their impact on $\mu$ by up to 65% when compared with the binned case [76].
We assess the performance with $5\cdot 10^{4}$ toys in Figure 4. The signal strength $\mu$ is reconstructed stably over the whole range of relevant $\mu_{\text{true}}$. We severely constrain $\nu_{\text{$\mathrm{t}\overline{\mathrm{t}}$}}$ and $\nu_{\text{jes}}$, reducing the impact of the corresponding uncertainties. The total training time for the model was 200 CPU core hours.
5.2 ibrahime: Contrastive Normalizing Flows for Uncertainty-Aware Parameter Estimation
The full description of the method can be found in the method paper [79]. The code used to train and evaluate the method is available at [80].
Motivation
A binary classifier can, in principle, estimate any model parameter $\Theta_{i}$ by learning a monotonic approximation of the likelihood ratio $r(\mathbf{x},\{\Theta_{i},\nu_{i}\},\{\Theta^{\prime}_{i},\nu^{\prime}_{i}\})\propto\frac{P(\mathbf{x}\!\mid\!\{\Theta_{i},\nu_{i}\})}{P(\mathbf{x}\!\mid\!\{\Theta^{\prime}_{i},\nu^{\prime}_{i}\})}$ [5], where $\mathbf{x}$ are the data features and $\nu_{i}$ are nuisance parameters. In practice, this classifier approach can be impractical; if the number of model parameters $k_{\Theta}$ or nuisance parameters $k_{\nu}$ is large, the dimensionality prevents sufficient sampling of parameter space for many choices of $\{\Theta_{i},\nu_{i}\}$. For the challenge, $\Theta\equiv\mu\propto f_{s}$, where $f_{s}$ is the signal fraction, and $\nu_{i}$ are the six HiggsML nuisance parameters. Given $\mu\propto f_{s}$ we can attempt to learn instead the likelihood ratio $r(\mathbf{x},\{\nu_{i}\},\{\nu^{\prime}_{i}\})\propto\frac{p_{s}(\mathbf{x}|\{\nu_{i}\})}{p_{b}(\mathbf{x}|\{\nu^{\prime}_{i}\})}$, where $p_{s}$ and $p_{b}$ are the signal and background distributions, by training on class labels and then determining $\mu$ with maximum likelihood estimation. To remedy the curse of dimensionality, we then replace the raw nuisance parameters $\nu_{i}$ with some discrimination functions $\Phi_{s,b}[\mathbf{x};\{\nu_{i}\}]$ such that $r(\mathbf{x},\{\nu_{i}\},\{\nu^{\prime}_{i}\})\propto\frac{p_{s}(\mathbf{x}|\Phi_{s}[\mathbf{x};\{\nu_{i}\}])}{p_{b}(\mathbf{x}|\Phi_{b}[\mathbf{x};\{\nu^{\prime}_{i}\}])}.$ If these discrimination functions are relatively insensitive to nuisance parameters and take very different values for $\mathbf{x}\sim p_{s}$ compared to $\mathbf{x}\sim p_{b}$, a classifier trained on these features will more accurately approximate the desired likelihood with less data. We argue that Contrastive Normalising Flows (CNFs) are especially suitable for these functions $\Phi_{s,b}[\mathbf{x};\{\nu_{i}\}]$.
Contrastive Normalising Flows (CNFs) A CNF is a normalising flow trained with a contrastive objective that simultaneously maximises the likelihood of one class and suppresses the likelihood of the other. Starting from the standard NF loss, and training on labelled data $\mathbf{x}_{s}\sim p_{s}$ and $\mathbf{x}_{b}\sim p_{b}$, we insert a term $c\,\log p_{\theta}^{(s)}(\mathbf{x}_{b})$ so that
$\begin{split}&\mathcal{L}_{s}=\frac{1}{|\mathcal{D}|}\sum_{\mathbf{x}_{s},\mathbf{x}_{b}\in\mathcal{D}}\Big\{-\log p_{\theta}^{(s)}(\mathbf{x}_{s})+c\,\log p_{\theta}^{(s)}(\mathbf{x}_{b})\Big\}\bigg\}\end{split}$ (5)
thereby causing the learned density $p_{\theta}^{(s)}$ to concentrate probability mass in regions characteristic of the signal and unlike background. CNFs have been used in anomaly detection settings [81]. We generalize with $c$ and develop a novel architecture and training procedure empirically required for accurate learning [79]. Exchanging the roles of $\mathbf{x}_{s}$ and $\mathbf{x}_{b}$ gives a loss function $\mathcal{L}_{b}$ and a learned function $p_{\theta}^{(b)}$ that concentrates in background regions. Transforming these probabilities as $\Phi_{s,b}(\mathbf{x})=p_{\theta}^{(s,b)}(\mathbf{x})/\bigl[1+p_{\theta}^{(s,b)}(\mathbf{x})\bigr]$ gives us our monotonic discrimination functions that retain the full shape of each class. Because the model learns a class distribution, not just a decision boundary, its scores are more stable under systematic shifts than those of a purely discriminative network. Tuning $c$ lets us trade off coverage versus stability under systematic shifts seen in Figure 5.
Figure 5: CNF distributions for various $c$ (left), DNN score histograms for signal varying the nuisance parameter $\alpha_{\rm tes}$ (center panel), the Neyman confidence belt (right)
The Method
The total training time is 10 GPU hours.
Step 1. Pre‑processing. Events are split into 1‑jet and 2‑jet categories (empirically, 0‑jet events hurt performance). We take the log of features which peak near zero and then standardise all features.
Step 2. CNF density learning. For each jet category we fit two CNFs $\bigl(p^{(s)}_{\theta,c},\,p^{(b)}_{\theta,c}\bigr)$ for $c\in\{0.5,\,2.0\}.$ $c>1$ sharpens signal‑rich regions and is empirically shift‑robust, while $c<1$ preserves coverage.
Step 3. DNN Classifier For any event $\mathbf{x}$ we compute $\Phi^{(s,b)}(\mathbf{x})=\frac{p^{(s,b)}_{\theta,c}(\mathbf{x})}{1+p^{(s,b)}_{\theta,c}(\mathbf{x})}$ for $c\in\{0.5,\,2.0\}$ yielding four CNF scores per jet category. Together with the primary and derived features, these are fed to a two‑headed DNN (shared trunk, jet‑specific heads) whose binary‑cross‑entropy loss is minimised on just 1,000 shifted mixtures uniformly sampling each $\nu_{i}$. We highlight the efficacy of CNF features with the relative invariance of the score histogram in Figure 5.
[…]
5.3 hzume: Decision‑Tree Aggregated Features and Hybrid Bin‑Classifier/Quantile‑Regressor
We build a two‑stage model composed of an Aggregation stage and an Estimation stage. Total training time is one CPU core hour.
Aggregation Stage: Feature Engineering : (i) For each event $(x_{ij})$ a decision tree estimates the class label $y_{ij}$ (signal vs. background), yielding a probability $p_{ij}$. From the set $\{p_{ij}\}$ we compute and aggregate mean, variance, skewness, kurtosis, and the empirical quantiles at levels 0255255. (ii) For each feature $x_{ij}$ its mean and variance across events, is fed into a second decision tree that predicts the Nuisance Parameters (e.g. TES, JES). These predictions are appended as additional features.
Estimation Stage: Two Models & Merging Strategy (i) A decision tree classifier partitions the interval $[0.1,3]$ into five equal‑width bins and predicts the bin containing $\mu$. The resulting probability is converted into the narrowest CI covering $68\,\%$ of the total probability. (ii) A quantile‑regression model directly predicts the lower and upper quantiles, providing an alternative CI for $\mu$.
Model Selection Rule. Empirically, the quantile regressor loses accuracy when $\mu$ is near the end‑points ($0.1$ or $3$). Therefore, we adopt the bin classifier in the edge regions and the quantile regressor in the central region to produce the final CI.
@@ 6 Conclusions and Outlook
We have prepared a dataset [65] (with relevant software [74]), challenge, and platform for developing and comparing machine learning methods that quantify uncertainties in addition to providing point estimates. With the growing size of datasets in high-energy physics, the sophistication of tools, and the precision requirements to explore new phenomena, uncertainty quantification will be an essential part of machine learning in the future. The two winning approaches, subsection 5.1 [76] and subsection 5.2 [79], show two alternative techniques on how the treatment of systematic uncertainties can be incorporated successfully in experimental analyses.
The two techniques have very similar performances, however their results are not very correlated which implies the optimum has not been reached yet. Beyond this specific metric, we expect that this unique large dataset equipped with a biasing script will be the basis of future studies, for example: (i) the precise parametrisation of density and density ratios over several order of magnitudes which is fundamental to precision physics (ii) development of morphing/Optimal Transport techniques to parameterise multidimensional non-parametric biases (iii) the same studies but with a focus on learning with a limited number of instances.